# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ Appendix A Large deviation estimates
We use a similar change of measure argument as in the proof of Proposition 10. Denote by $I$ the large deviation rate function for the $\mathbb{R}^{d}$-valued random variable ${\bm{\xi}}$. For $\widehat{\mathbf{c}}_{2}:=\nabla I({\mathbf{z}}/(Ck))$, we define the tilted measure $\mathbb{Q}$ by $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}(\mathbf{x}):=e^{\widehat{\mathbf{c}}_{2}\cdot\mathbf{x}-\Lambda(\widehat{\mathbf{c}}_{2})}$, where we recall $\Lambda=\log\phi_{\bm{\xi}}$ is the log-moment generating function. Under $\mathbb{Q}$, each ${\bm{\xi}}_{i}$ is i.i.d. with mean ${\mathbf{z}}/(Ck)$. It follows that
$\mathbb{P}((X,\mathbf{Y})\in B_{{\mathbf{z}}})$ $=e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})}\mathbb{E}_{\mathbb{Q}}[e^{-\widehat{\mathbf{c}}_{2}\cdot(X,\mathbf{Y})}{\mathbbm{1}}_{\{(X,\mathbf{Y})\in B_{{\mathbf{z}}}\}}]$
$\gg e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})-\widehat{\mathbf{c}}_{2}\cdot{\mathbf{z}}}\mathbb{Q}\left(\left(\frac{X}{\sqrt{Ck}},\frac{\mathbf{Y}}{\sqrt{Ck}}\right)\in\frac{B_{{\mathbf{z}}}}{\sqrt{Ck}}\right)$
$=e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})-\widehat{\mathbf{c}}_{2}\cdot{\mathbf{z}}}\mathbb{Q}\left(\left(\frac{\widetilde{X}}{\sqrt{Ck}},\frac{\widetilde{\mathbf{Y}}}{\sqrt{Ck}}\right)\in\frac{B_{\mathbf{0}}}{\sqrt{Ck}}\right),$ (67)
where each $(\widetilde{X_{i}},\widetilde{\mathbf{Y}}_{i})=(X_{i},\mathbf{Y}_{i})-{\mathbf{z}}/(Ck)$ is a centered random variable. Using definition of $\widehat{\mathbf{c}}_{2}$, we have
@@
In addition, Lemma 24 yields
$\mathbb{Q}\left(\left(\frac{\widetilde{X}}{\sqrt{Ck}},\frac{\widetilde{\mathbf{Y}}}{\sqrt{Ck}}\right)\in\frac{B_{\mathbf{0}}}{\sqrt{Ck}}\right)\gg(Ck)^{-d/2}.$ (69)
@@
$\mathbb{P}((X,\mathbf{Y})\in B_{{\mathbf{z}}})\gg e^{-CkI({\mathbf{z}}/(Ck))}(Ck)^{-d/2}.$
Recalling that $I$ is convex, $C^{1}$ on its domain, and $I(\mathbf{0})=\nabla I(\mathbf{0})=0$ (by Jensen’s inequality, $\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ is minimized at $\bm{\lambda}=\mathbf{0}$ when ${\bm{\xi}}$ is centered), we have
@@
The Bahadur-Rao theorem yields that uniformly in $c\in[-C\log n,C\log n]$,
$\mathbb{P}(X\geqslant x+c)\asymp n^{-1/2}e^{-nI(x/n)-cI^{\prime}(x/n)}.$ (70)
Mimicing (65), define the probability measure $\mathbb{Q}$ by
@@
It follows that $\mathbb{E}_{\mathbb{Q}}[{\bm{\xi}}]=(x/n,\mathbf{0})$. That is, the random variable $(\widetilde{X}_{i},\mathbf{Y}_{i})$ is then centered under $\mathbb{Q}$ where $\widetilde{X}_{i}=X_{i}-x/n$. Define $\widetilde{X}:=\widetilde{X}_{1}+\dots+\widetilde{X}_{n}$. It follows that
$\hskip 14.22636pt\mathbb{P}(X\geqslant x+c,\,\mathbf{Y}\in B_{\mathbf{0}}(1))$
$=e^{n\Lambda(\mathbf{c}_{2})}\mathbb{E}_{\mathbb{Q}}\left[e^{-\mathbf{c}_{2}\cdot(X,\mathbf{Y})}{\mathbbm{1}}_{\{X\geqslant x+c,\,\mathbf{Y}\in B_{\mathbf{0}}(1)\}}\right]$
$\asymp e^{n\Lambda(\mathbf{c}_{2})-\mathbf{c}_{2}\cdot(x,\mathbf{0})}\sum_{j=c+1}^{\infty}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\widetilde{X}\in[j-1,j),\,\mathbf{Y}\in B_{\mathbf{0}}(1)\right)$
$=e^{-nI(x/n,\mathbf{0})}\sum_{j=c+1}^{\infty}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\frac{(\widetilde{X},\mathbf{Y})}{\sqrt{n}}\in\Big{[}\frac{j-1}{\sqrt{n}},\frac{j}{\sqrt{n}}\Big{)}\times B_{\mathbf{0}}\Big{(}\frac{1}{\sqrt{n}}\Big{)}\right)$
@@
$\mathbb{P}(X\in[x+c,x+c+a(n)])\asymp n^{-1/2}e^{-nI(x/n)-cI^{\prime}(x/n)}$
@@
$\hskip 14.22636pt\mathbb{P}(\mathbf{Y}\in B_{\mathbf{y}}(1),\leavevmode\nobreak\ X\in[x+c,x+c+a(n)])$
$\asymp e^{n\Lambda(\mathbf{c}_{2})-\mathbf{c}_{2}\cdot(x,\mathbf{y})}\sum_{j=c+1}^{c+a(n)}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\frac{(\widetilde{X},\mathbf{Y})}{\sqrt{n}}\in\Big{[}\frac{j-1}{\sqrt{n}},\frac{j}{\sqrt{n}}\Big{)}\times B_{\frac{\mathbf{y}}{\sqrt{n}}}\Big{(}\frac{1}{\sqrt{n}}\Big{)}\right)$

### v2

@@ 5 Concluding remarks
We use a similar change of measure argument as in the proof of Proposition 10. Denote by $I$ the large deviation rate function for the $\mathbb{R}^{d}$-valued random variable $\bm{\xi}$} (defined in (2)). For $\widehat{\mathbf{c}}_{2}:=\nabla I({\mathbf{z}}/(Ck))$, we define the tilted measure $\mathbb{Q}$ by $\mathrm{d}\mathbb{Q}/\mathrm{d}\mathbb{P}(\mathbf{x}):=e^{\widehat{\mathbf{c}}_{2}\cdot\mathbf{x}-\Lambda(\widehat{\mathbf{c}}_{2})}$, where we recall $\Lambda=\log\phi_{\bm{\xi}}$ is the log-moment generating function. Under $\mathbb{Q}$, each ${\bm{\xi}}_{i}$ is i.i.d. with mean ${\mathbf{z}}/(Ck)$. It follows that
$\mathbb{P}((F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}})$ $=e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})}\mathbb{E}_{\mathbb{Q}}[e^{-\widehat{\mathbf{c}}_{2}\cdot(F_{Ck},\mathbf{J}_{Ck})}{\mathbbm{1}}_{\{(F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}}\}}]$
$\gg e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})-\widehat{\mathbf{c}}_{2}\cdot{\mathbf{z}}}\mathbb{Q}\left(\left(\frac{F_{Ck}}{\sqrt{Ck}},\frac{\mathbf{J}_{Ck}}{\sqrt{Ck}}\right)\in\frac{B_{{\mathbf{z}}}}{\sqrt{Ck}}\right)$
$=e^{Ck\Lambda(\widehat{\mathbf{c}}_{2})-\widehat{\mathbf{c}}_{2}\cdot{\mathbf{z}}}\mathbb{Q}\left(\left(\frac{\widetilde{F}_{Ck}}{\sqrt{Ck}},\frac{\widetilde{\mathbf{J}}_{Ck}}{\sqrt{Ck}}\right)\in\frac{B_{\mathbf{0}}}{\sqrt{Ck}}\right),$ (68)
where each $(\widetilde{X_{i}},\widetilde{\mathbf{Y}}_{i})=(X_{i},\mathbf{Y}_{i})-{\mathbf{z}}/(Ck)$ is a centered random variable and $(\widetilde{F}_{Ck},\widetilde{\mathbf{J}}_{Ck})=\sum_{j=1}^{Ck}(\widetilde{X_{i}},\widetilde{\mathbf{Y}}_{i})$. Using definition of $\widehat{\mathbf{c}}_{2}$, we have
@@
In addition, Lemma 24 yields [^15] Lemma 24 is stated for radially symmetric distributions, but the part that derives (64) from (63) does not depend on the symmetry and applies to general non-lattice distributions. Note that the non-lattice condition (A5) is preserved under exponential tilts.
$\mathbb{Q}\left(\left(\frac{\widetilde{F}_{Ck}}{\sqrt{Ck}},\frac{\widetilde{\mathbf{J}}_{Ck}}{\sqrt{Ck}}\right)\in\frac{B_{\mathbf{0}}}{\sqrt{Ck}}\right)\gg(Ck)^{-d/2}.$ (70)
@@
$\mathbb{P}((F_{Ck},\mathbf{J}_{Ck})\in B_{{\mathbf{z}}})\gg e^{-CkI({\mathbf{z}}/(Ck))}(Ck)^{-d/2}.$
Recalling that $I$ is convex, $C^{1}$ on its domain, and $I(\mathbf{0})=\nabla I(\mathbf{0})=0$ (by Jensen’s inequality, $\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ is minimized at $\bm{\lambda}=\mathbf{0}$ when $\bm{\xi}$} is centered), we have uniformly for $\mathbf{z}$} with $\left\lVert{\mathbf{z}}\right\rVert\leqslant C_{1}k$,
@@
The Bahadur–Rao theorem (Theorem 3.7.4 of [dembo2009large]) yields that uniformly in $c\in[-C\log n,C\log n]$,
$\mathbb{P}(F_{n}\geqslant x+c)\asymp n^{-1/2}e^{-nI(x/n)-cI^{\prime}(x/n)}.$ (71)
Mimicking (66), define the probability measure $\mathbb{Q}$ by
@@
It follows that $\mathbb{E}_{\mathbb{Q}}[{\bm{\xi}}]=(x/n,\mathbf{0})$. That is, the random variable $(\widetilde{X}_{i},\mathbf{Y}_{i})$ is then centered under $\mathbb{Q}$ where $\widetilde{X}_{i}=X_{i}-x/n$. Define $\widetilde{F}_{n}:=\widetilde{X}_{1}+\dots+\widetilde{X}_{n}$. It follows that
$\hskip 14.22636pt\mathbb{P}(F_{n}\geqslant x+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(1))$
$=e^{n\Lambda(\mathbf{c}_{2})}\mathbb{E}_{\mathbb{Q}}\left[e^{-\mathbf{c}_{2}\cdot(F_{n},\mathbf{J}_{n})}{\mathbbm{1}}_{\{F_{n}\geqslant x+c,\,\mathbf{J}_{n}\in B_{\mathbf{0}}(1)\}}\right]$
$\asymp e^{n\Lambda(\mathbf{c}_{2})-\mathbf{c}_{2}\cdot(x,\mathbf{0})}\sum_{j=c+1}^{\infty}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\widetilde{F}_{n}\in[j-1,j),\,\mathbf{J}_{n}\in B_{\mathbf{0}}(1)\right)$
$=e^{-nI(x/n,\mathbf{0})}\sum_{j=c+1}^{\infty}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1}{\sqrt{n}},\frac{j}{\sqrt{n}}\Big)\times B_{\mathbf{0}}\Big(\frac{1}{\sqrt{n}}\Big)\right)$
@@
$\mathbb{P}(F_{n}\in[x+c,x+c+a(n)])\asymp n^{-1/2}e^{-nI(x/n)-cI^{\prime}(x/n)}$
@@
$\hskip 14.22636pt\mathbb{P}(\mathbf{J}_{n}\in B_{\mathbf{y}}(1),\penalty 10000\ F_{n}\in[x+c,x+c+a(n)])$
$\asymp e^{n\Lambda(\mathbf{c}_{2})-\mathbf{c}_{2}\cdot(x,\mathbf{y})}\sum_{j=c+1}^{c+a(n)}e^{-\mathbf{c}_{2}\cdot(j,\mathbf{0})}\mathbb{Q}\left(\frac{(\widetilde{F}_{n},\mathbf{J}_{n})}{\sqrt{n}}\in\Big[\frac{j-1}{\sqrt{n}},\frac{j}{\sqrt{n}}\Big)\times B_{\frac{\mathbf{y}}{\sqrt{n}}}\Big(\frac{1}{\sqrt{n}}\Big)\right)$